# VDGD: Mitigating LVLM Hallucinations in Cognitive Prompts by Bridging the Visual Perception Gap

### v1

@@ H.6 Examples of Rephrased Prompts
Table 9: Performance comparison of various LVLMs on MathVista benchmark. This is an extension of results shown in Figure 7.
Model Methodology Helpfulness Clarity Factuality Depth Engagement
LLaVA-v1 Vanilla-greedy 3.12 4.41 2.44 2.19 1.23
Vanilla-sampling 3.14 4.42 2.45 2.20 1.26
VDD 2.98 4.10 2.20 2.07 1.16
OPERA 3.20 4.48 2.50 2.22 1.28
Woodpecker 3.21 4.49 2.46 2.21 1.25
LRV 3.25 4.51 2.51 2.23 1.29
LURE 3.20 4.47 2.49 2.20 1.28
VDGD (ours) 3.42 4.59 2.78 2.30 1.30
LLaVA-1.5 Vanilla-greedy 3.25 4.41 2.66 2.61 1.19
Vanilla-sampling 3.24 4.40 2.65 2.58 1.10
VDD 2.82 4.10 2.44 2.39 1.20
OPERA 3.22 4.40 2.49 2.58 1.18
Woodpecker 3.26 4.42 2.46 2.64 1.22
LRV 3.25 4.38 2.52 2.59 1.24
LURE 3.24 4.35 2.51 2.60 1.21
VDGD (ours) 3.47 4.56 2.84 2.51 1.80
LLaVA-1.6 Vanilla-greedy 3.43 4.41 3.13 3.31 1.35
Vanilla-sampling 3.40 4.38 3.10 3.27 1.30
VDD 3.26 4.19 2.96 3.13 1.20
OPERA 3.49 4.48 3.20 3.38 1.40
Woodpecker 3.45 4.45 3.21 3.35 1.42
LRV 3.42 4.40 3.18 3.34 1.39
LURE 3.38 4.36 3.18 3.28 1.36
VDGD (ours) 3.64 4.71 3.49 3.42 1.45
mPLUG-Owl2 Vanilla-greedy 3.23 4.42 2.86 2.48 1.25
Vanilla-sampling 3.24 4.47 2.84 2.50 1.24
VDD 3.07 4.21 2.67 2.40 1.17
OPERA 3.29 4.52 2.92 2.58 1.27
Woodpecker 3.27 4.50 2.90 2.52 1.25
LRV 3.23 4.49 2.88 2.50 1.24
LURE 3.26 4.51 2.91 2.51 1.28
VDGD (ours) 3.34 4.64 3.15 2.70 1.32
InternLM-X Vanilla-greedy 3.28 4.52 3.35 2.59 1.37
Vanilla-sampling 3.27 4.53 3.33 2.57 1.35
VDD 3.08 4.24 3.10 2.44 1.30
OPERA 3.37 4.65 3.44 2.65 1.43
Woodpecker 3.38 4.63 3.45 2.66 1.44
LRV 3.40 4.69 3.47 2.70 1.46
LURE 3.36 4.64 3.42 2.68 1.42
VDGD (ours) 3.54 4.78 3.68 2.79 1.49
CogVLM Vanilla-greedy 3.58 4.51 3.35 2.76 1.38
Vanilla-sampling 3.59 4.50 3.37 2.78 1.41
VDD 3.36 4.32 3.21 2.58 1.30
OPERA 3.68 4.60 3.42 2.84 1.45
Woodpecker 3.70 4.58 3.45 2.85 1.42
LRV 3.65 4.57 3.40 2.82 1.40
LURE 3.72 4.62 3.46 2.83 1.41
VDGD (ours) 3.89 4.82 3.59 2.89 1.49
GPT-4-Turbo - 4.15 4.79 3.89 3.25 1.67
@@
H.7 Compute Infrastructure
Appendix I Additional Results
I.1 Comparison of various LVLMs on rephrased prompts
I.2 Comparison of various LVLMs on image description
I.3 Comparison of post-truncation probability statistics on AMBER and MMMU benchmarks
I.4 Comparison of LVLMs on different capabilities on the proposed VaLLu benchmark
Table 3 provides comparison of LVLMs on different capabilities on the proposed VaLLu benchmark.
I.5 Comparison of VaLLu with Mitigation Baselines
I.6 Comparison of MMMU with Mitigation Baselines
I.7 Comparison of MathVista with Mitigation Baselines
I.8 Comparison of Oven with Mitigation Baselines
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, the main claims made in the abstract and introduction accurately reflect the paper’s contribution and are true to the best of our knowledge.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: The paper’s last section i.e 7 Conclusion, Limitations and Future Work discusses the limitation of our work.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: This paper does not make any theoretical assumptions and does not describe any theoretical proofs.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: To the best of our knowledge we described our experimental methods, algorithms and evaluation prompts in the paper in detail.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: We have released the code for all the experiments, algorithm and our novel contributions in an anonymous git reopsitory as described in 7. We have also uploaded the benchmark proposed in the paper under Supplementary Materials.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We have included all the evaluation prompts and hyperparameters used for testing. The paper does not provide any training methodologies and as such does not specify any training details.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: We have provided error bars for the results in the main paper. Due to the expensive nature of evaluation, we provided error bars only for the main results in the paper.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: We have details about all the compute resources we have used in section H.7.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: To the best of our knowledge the paper adheres to all the ethical guidelines provided in the above link.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: We have described the broader/societal impacts of the paper in section C.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [No]
  44. Justification: We do not release any new data or models. All data and models used are from existing resources. The VaLLu benchmark is an amalgamation of existing benchmarks. We use licenses and release the dataset under licenses issued by the original benchmark.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: All the licenses for models, datasets and baselines used in the paper are provided in appendix sections F, E and G.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We introduced a new benchmark in this paper. The benchmark data and metadata file have been uploaded in the supplementary folder in an anonymized zip file.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [No]
  56. Justification: We did not use crowdsourcing and research with human subjects.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [Yes]
  60. Justification: We obtained an IRB approval to filter out noisy examples in the datasets we used as described in section H.3.

### v2

@@ Appendix I Model Details
Table 11: Performance comparison of various LVLMs on MathVista benchmark (only questions tagged with open-ended generation). This is an extension of results shown in Figure 7.
[…]
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.12 4.41 2.44 2.19 1.23 Vanilla-sampling 3.14 4.42 2.45 2.20 1.26 VDD 2.98 4.10 2.20 2.07 1.16 OPERA 3.20 4.48 2.50 2.22 1.28 Woodpecker 3.21 4.49 2.46 2.21 1.25 LRV 3.25 4.51 2.51 2.23 1.29 LURE 3.20 4.47 2.49 2.20 1.28 VDGD (ours) 3.42 4.59 2.78 2.30 1.30 LLaVA-1.5 Vanilla-greedy 3.25 4.41 2.66 2.61 1.19 Vanilla-sampling 3.24 4.40 2.65 2.58 1.10 VDD 2.82 4.10 2.44 2.39 1.20 OPERA 3.22 4.40 2.49 2.58 1.18 Woodpecker 3.26 4.42 2.46 2.64 1.22 LRV 3.25 4.38 2.52 2.59 1.24 LURE 3.24 4.35 2.51 2.60 1.21 VDGD (ours) 3.47 4.56 2.84 2.51 1.80 LLaVA-1.6 Vanilla-greedy 3.43 4.41 3.13 3.31 1.35 Vanilla-sampling 3.40 4.38 3.10 3.27 1.30 VDD 3.26 4.19 2.96 3.13 1.20 OPERA 3.49 4.48 3.20 3.38 1.40 Woodpecker 3.45 4.45 3.21 3.35 1.42 LRV 3.42 4.40 3.18 3.34 1.39 LURE 3.38 4.36 3.18 3.28 1.36 VDGD (ours) 3.64 4.71 3.49 3.42 1.45 mPLUG-Owl2 Vanilla-greedy 3.23 4.42 2.86 2.48 1.25 Vanilla-sampling 3.24 4.47 2.84 2.50 1.24 VDD 3.07 4.21 2.67 2.40 1.17 OPERA 3.29 4.52 2.92 2.58 1.27 Woodpecker 3.27 4.50 2.90 2.52 1.25 LRV 3.23 4.49 2.88 2.50 1.24 LURE 3.26 4.51 2.91 2.51 1.28 VDGD (ours) 3.34 4.64 3.15 2.70 1.32 InternLM-X Vanilla-greedy 3.28 4.52 3.35 2.59 1.37 Vanilla-sampling 3.27 4.53 3.33 2.57 1.35 VDD 3.08 4.24 3.10 2.44 1.30 OPERA 3.37 4.65 3.44 2.65 1.43 Woodpecker 3.38 4.63 3.45 2.66 1.44 LRV 3.40 4.69 3.47 2.70 1.46 LURE 3.36 4.64 3.42 2.68 1.42 VDGD (ours) 3.54 4.78 3.68 2.79 1.49 CogVLM Vanilla-greedy 3.58 4.51 3.35 2.76 1.38 Vanilla-sampling 3.59 4.50 3.37 2.78 1.41 VDD 3.36 4.32 3.21 2.58 1.30 OPERA 3.68 4.60 3.42 2.84 1.45 Woodpecker 3.70 4.58 3.45 2.85 1.42 LRV 3.65 4.57 3.40 2.82 1.40 LURE 3.72 4.62 3.46 2.83 1.41 VDGD (ours) 3.89 4.82 3.59 2.89 1.49 GPT-4-Turbo - 4.15 4.79 3.89 3.25 1.67